OpenStudy (anonymous):

what number is 133 1/3 of 36?

OpenStudy (texaschic101):

x = 133 1/3 (36) x = 400/3 (36) x = 14400/3 = 4800 check... 4800 = 400/3 * 36 4800 = 4800 (correct)

OpenStudy (anonymous):

thanks you :)!